OpenStudy (anonymous):

(2/x)^2-4=7 Pls solve

OpenStudy (zepp):

\[(\frac{2}{x})^2-4=7\] Take off the 4\[(\frac{2}{x})^2-4+4=7+4\]\[(\frac{2}{x})^2=11\] Take the square root on both side \[(\frac{2}{x})=\sqrt{11}\] Multiply by one half.. \[(\frac{2}{\sqrt{11}})=x\]

OpenStudy (anonymous):

pls tell the a,b and c

OpenStudy (zepp):

Please tell me what is A, B and C from what you learnt?